OpenStudy (anonymous):

an object is dropped off a cliff. how far will the object fall during the next 4s?

OpenStudy (aaronq):

it's in constant acceleration since the only force on it is gravity \(\Delta x=v_i\Delta t+\dfrac{1}{2}a\Delta t^2\) \(\Delta x=0(4s)+\dfrac{1}{2}g(4s)^2\) \(\Delta x=\dfrac{1}{2}(-9.8~m/s^2)(4s)^2\) \(\Delta x=\dfrac{1}{2}(-9.8~m/s^2)(4s)^2=-78.4 ~m\) it's negative because its in relation to the y axis.